OpenStudy (anonymous):

Michelle made a bead necklace that was 14 inches long. The length of the necklace, in inches, is equal to 4 added to one-fifth the number of beads, b. 1/5b+4=14 How many beads did Michelle use for her necklace? A.15 B.50 C.500 D.25

OpenStudy (danjs):

hint : multiply everything by 5 first

Nnesha (nnesha):

or you can replace answer choices by variable b then solve left side if you get both side equal then that is your aanswer

OpenStudy (anonymous):

it is 50

OpenStudy (danjs):

\[\frac{ 5*1 }{ 5 }*b + 5*4 = 5*14\] \[b + 20 = 70\]

OpenStudy (danjs):

yep

OpenStudy (anonymous):

thxs

Nnesha (nnesha):

lets try  both side equal then yes that is your right answer